OpenStudy (satan312):

i need help with a slope y+7=-3/4(x-4)

OpenStudy (puppylover784):

Hi ZAY

OpenStudy (puppylover784):

x in (-oo:+oo) Y-7 = (-3/4)*(x-4) // - (-3/4)*(x-4) Y-((-3/4)*(x-4))-7 = 0 (3/4)*(x-4)+Y-7 = 0 3/4*(x-4)+Y-7 = 0 3/4*(x-4)+Y-7 = 0 Y+3/4*x-7-3 = 0 Y+3/4*x-10 = 0 Y+3/4*x-10 = 0 Y+3/4*x-10 = 0 // - Y-10 3/4*x = -(Y-10) // : 3/4 x = (-(Y-10))/3/4 x = 4/3*(10-Y) x = 4/3*(10-Y)

OpenStudy (satan312):

thanks

OpenStudy (satan312):

can you help me with a few more

OpenStudy (puppylover784):

Sure

OpenStudy (satan312):

(-3,4;m=7

OpenStudy (puppylover784):

y=7x+25

OpenStudy (satan312):

ok thank you